# The Uniform Mordell–Lang Conjecture

### v3

@@ Abstract.
The Mordell–Lang conjecture for abelian varieties states that the intersection of an algebraic subvariety $X$ with a subgroup of finite rank is contained in a finite union of cosets contained in $X$. In this article, we prove a uniform version of this conjecture, meaning that that the number of cosets necessary does not depend on the ambient abelian variety. To achieve this, we prove a general gap principle on algebraic points that extends the gap principle for curves embedded into their Jacobians, previously obtained by Dimitrov–Gao–Habegger and Kühne. Our new gap principle also implies the full uniform Bogomolov conjecture in abelian varieties.
@@ 1. Introduction
Throughout this article, $F$ is an algebraically closed field of characteristic 00; particularly relevant cases are $F=\overline{\mathbb{Q}}$ and $F=\mathbb{C}$. Furthermore, we let $A$ be an abelian variety defined over $F$ and consider an ample line bundle $L$ on $A$. The translates of abelian subvarieties by closed points in $A$ are called cosets.
@@
The novelity here is the complete removal of this very dependence on the ambient abelian variety $A$, confirming a folklore expectation that can be found for example in [DP07, Conj. 1.8]. It is known under the name of uniform Mordell–Lang conjecture because the number $N$ in (1.1) must depend on $g$, $d$ and $\rho$. For curves embedded in their Jacobian, it dates back to a question of Mazur [Maz86, top of p. 234], which has been answered affirmatively by work of Dimitrov, the first-named author, Habegger and the third-named author [DGH21, Küh21]. Readers may profit from the survey of the first-named author [Gao21] for an overview of these previous works and their implications for rational points on algebraic curves. Let us remark that before these works, the only uniform results of Mordell-Lang type were obtained by David and Philippon [DP07, Théorème 1.13] for subvarieties of self-products of an elliptic curve. It should be also noted that they give a completely explicit constant in this special case. In this regard, it is interesting to ask whether the present arguments can yield explicit upper bounds on the number of cosets similar to those of David and Philippon [DP07] or if substantial new ideas are necessary.
@@ 1.1. The Ueno locus and first reductions
(1.2) #$\#X^{\circ}(F)\cap\Gamma\leq c(\dim A,\deg_{L}X)^{1+\mathrm{rk}\Gamma}.$
Besides reducing to Theorem 1.1, we also use a specialization argument of Masser [Mas89] to assume $F=\overline{\mathbb{Q}}$ in our main arguments. This is essential for most of our arithmetically flavored techniques revolving around the theory of heights. Although Theorem 1.1 could itself be understood as a purely geometric assertion (e.g. for $F=\mathbb{C}$), we heavily rely on arithmetic tools throughout our proof.
@@ 1.2. A generalized gap principle
Our main contribution is a uniform count of small points, and we achieve this by establishing another kind of gap principle for algebraic points. In the case of curves in their Jacobians, preceding work [DGH21, Küh21] has been subsumed under such a New Gap Principle [Gao21, Thm. 4.1]. In this article, we take the same perspective and generalize it as follows. We say that a subvariety $X\subseteq A$ generates $A$ if the smallest abelian subvariety containing $X-X$ is $A$. Furthermore, we let $\hat{h}_{L\otimes[-1]^{*}L}$ denote the Néron–Tate height on $A(\overline{\mathbb{Q}})$ associated with the symmetric ample line bundle $L\otimes[-1]^{*}L$.
@@
There exist positive constants $c_{1}=c_{1}(\dim A,\deg_{L}X)$ and $c_{2}=c_{2}(\dim A,\deg_{L}X)$ with the following property: For any irreducible closed subvariety $X\subseteq A$ that generates $A$, the set
@@
is contained in some Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}(X^{\prime})<c_{2}$.
@@
We remark that Theorem 1.2 implies a uniform version of the Bogomolov conjecture, generalizing [Küh21, Thm.3] in the case of curves embedded in their Jacobians. Since this result is of independent interest, we state it here explicitly as well.
@@
There exist positive constants $c_{3}=c_{3}(\dim A,\deg_{L}X)$ and $c_{4}=c_{4}(\dim A,\deg_{L}X)$ with the following property: For each irreducible subvariety $X$ of $A$, we have
(1.4) #$\#\left\{P\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes[-1]^{*}L}(P)\leq c_{3}\right\}<c_{4}.$
@@
An analogue of Theorem 1.3 for curves in algebraic tori was proven by Bombieri–Zannier [BZ95] (compare also [DP99, AD06]). For the case of abelian varieties considered here, Theorem 1.3 was known in some selected cases [DP07, DKY20, Küh21]. Let us note that all these results except for [Küh21] have rather explicit constants, in contrast to our general Theorem 1.3 here.
After the first version of the current article appeared as a preprint, Yuan [Yua21] gave another proof of Theorem 1.2 in the case of curves embedded into their Jacobians previously considered in [DGH21, Küh21]. His approach has the advantage to work in the function field case (in positive characteristic) as well. It relies on previous work of Zhang, Cinkir, and de Jong [Zha93, Zha10, Cin11, dJ18] for lower bounds on the self-intersection numbers of the admissible canonical bundles of curves over global fields. Little is known in this direction for subvarieties of higher dimension and it seems very hard to generalize Yuan’s proof to the setting considered here. He also uses the new theory of Yuan–Zhang [YZ21] on adelic line bundles over quasi-projective varieties.
@@ 1.3. Ideas of the proof
    the height inequality of [DGH21, Thm. 1.6 and B.1], which allows a comparison of the Néron–Tate height and the height on the base variety, as well as
@@
More recently, Yuan and Zhang have reproven both these results using their general theory of adelic line bundles over quasi-projective varieties [YZ21].
@@
To prove uniform bounds as in Theorem 1.1, it is important to work with only finitely many families at all times. While this is automatic in the setting for curves embedded into their Jacobians (by an induction on the dimension of the subvariety), we have to carefully handle the invariants involved, in particular the polarization type. Even if one is only interested in the case of principal polarization, our inductive proof generally invokes abelian variety of any polarization degree. We use various techniques to overcome these problems, including a classical result of Mumford and the Poincaré biextension on the universal abelian variety.
@@ Acknowledgements
The authors would like to thank Marc Hindry for the argument for Lemma 2.5. The authors would like to thank Gabriel Dill for his valuable comments on a previous version of the manuscript, especially for a correction of the argument for Lemma 3.1, the reference [Dil22, Lemma 2.4], discussions on the specialization argument and on the optimality of our formulation of Theorem 1.1. We thank Dan Abramovich and Philipp Habegger for comments on a draft of this paper and Abbey Bourdon for a relevant discussion on the consequence of our main result on rational points. Ziyang Gao received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n 945714). Tangli Ge received funding from NSF grant DMS-1759514 and DMS-2100548. Lars Kühne received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101027237.
@@ 2. Preliminaries on abelian varieties
In this section, we include some basic results on abelian varieties. Throughout the section, let $A$ be an abelian variety defined over $\overline{\mathbb{Q}}$ of dimension $g$.
@@ 2.1. Polarizations
Let $\mathrm{Pic}(A)$ be the Picard group of $A$ and $\mathrm{Pic}^{0}(A)$ be the connected component of the identity. For $L\in\mathrm{Pic}(A)$, denote its Chern class by $c_{1}(L)\in H^{2}(A(\mathbb{C}),\mathbb{Z})$. Let $A^{\vee}$ be the dual abelian variety of $A$. Then $\mathrm{Pic}^{0}(A)=A^{\vee}(\overline{\mathbb{Q}})$.
@@
Convention. We will use the following convention. By a polarized abelian variety, we mean a pair $(A,L)$ where $A$ is an abelian variety and $L$ is an ample line bundle. We say that $(A,L)$ is defined over $\overline{\mathbb{Q}}$ if both $A$ and $L$ are defined over $\overline{\mathbb{Q}}$.
The following Lemma is [Deb05, Theorem 6.10], which says that the polarization is uniquely determined by the Chern class of the line bundle.
@@
Next we define the polarization type. Let $(A,L)$ be a polarized abelian variety.
@@
    let $f\colon A^{\prime}\rightarrow A$ be an isogeny, then $\dim H^{0}(A^{\prime},f^{*}L)=\deg(f)\dim H^{0}(A,L)$.
@@
    there exist an abelian variety $A_{0}$, an ample line bundle $L_{0}$ on $A_{0}$ defining a principal polarization, and an isogeny $u_{0}\colon A\rightarrow A_{0}$ such that $L\cong u_{0}^{*}L_{0}$; moreover, $\deg(u_{0})=\deg_{L}(A)/g!$.
@@
For (i), (ii) and (iii), it suffices to prove the assertions over $\mathbb{C}$. Then (i) is [BL04, Corollary 3.2.8], (ii) is just the Riemann–Roch theorem [BL04, Section 3.6] and (iii) is [Deb05, Cororollary 6 to Proposition 6.12]. (iv) is [Mum74, pp. 216, Corollary 1 and its proof]. ∎
@@
We have $c_{1}(L^{\prime}\otimes L^{\otimes-1})=c_{1}(L^{\prime})-c_{1}(L)=0$. So $L^{\prime}\otimes L^{\otimes-1}\in\mathrm{Pic}^{0}(A)=A^{\vee}(\overline{\mathbb{Q}})$. The morphism of abelian varieties $\phi_{L}\colon A\rightarrow A^{\vee}$, $a\mapsto t_{a}^{*}L\otimes L^{\otimes-1}$, is an isogeny because $L$ is ample. So there exists $a\in A(\overline{\mathbb{Q}})$ such that $\phi_{L}(a)=L^{\prime}\otimes L^{\otimes-1}$. Thus $L^{\prime}\cong t_{a}^{*}L$. Therefore for each $x\in A(\overline{\mathbb{Q}})$, we have
@@
Take $x\in A(\overline{\mathbb{Q}})_{\mathrm{tor}}$, then $\hat{h}_{L^{\prime}}(x)=\hat{h}_{L}(x)=0$ for the symmetric ample line bundles $L$ and $L^{\prime}$. So (2.3) yields $0=\hat{h}_{L}(a)$. But then $a\in A(\overline{\mathbb{Q}})_{\mathrm{tor}}$ since $L$ is symmetric ample, and hence (2.3) yields $\hat{h}_{L^{\prime}}(x)=\hat{h}_{L}(x)$ for all $x\in A(\overline{\mathbb{Q}})$. ∎
@@ 2.2. Degree estimates
$=\sum_{i=0}^{d+d^{\prime}}\binom{d+d^{\prime}}{i}(c_{1}(p_{1}^{*}L)^{i}\cap Y)\times(c_{1}(p_{2}^{*}L)^{d+d^{\prime}-i}\cap Y^{\prime}).$
For reasons of dimension, the only non-vanishing term in this sum is for $i=d$. Therefore,
@@
$\deg_{\alpha^{*}L^{\boxtimes 2}}(Y\times Y^{\prime})=2^{d+d^{\prime}}\binom{d+d^{\prime}}{d}\deg_{L}Y\cdot\deg_{L}Y^{\prime}.$
@@
Let $X$ be an irreducible subvariety of $A$ and let $A^{\prime}$ denote the abelian subvariety generated by $X-X$. Then $\deg_{L}A^{\prime}\ll_{g}\deg_{L}(X)^{2g}$.
@@
$A^{\prime}=\underbrace{(X-X)+(X-X)+\cdots+(X-X)}_{\text{$g$ copies}}.$
@@
(2.7) $\deg_{L}(\underbrace{(X-X)+\cdots+(X-X)}_{\text{$k$ copies}})\leq 8^{2^{k}r}\deg_{L}(X)^{2g}$
@@ 3. Hilbert schemes and non-degeneracy
Let $r\geq 1$ and $d\geq 1$ be two integers. In this section, we work over $\overline{\mathbb{Q}}$, i.e. all objects and morphisms are defined over $\overline{\mathbb{Q}}$ unless said otherwise.
The goal of this section is to introduce the restricted Hilbert schemes (3.6) and prove a non-degeneracy result, Proposition 3.4. This is a main new ingredient to prove the Uniform Mordell–Lang Conjecture for higher dimensional subvarieties of abelian varieties (compared to the case of curves).
@@
3.1. Abelian variety

### v4

@@ Abstract.
The Mordell–Lang conjecture for abelian varieties states that the intersection of an algebraic subvariety $X$ with a subgroup of finite rank is contained in a finite union of cosets contained in $X$. In this article, we prove a uniform version of this conjecture, meaning that the number of cosets necessary depend only on the dimension of the ambient variety and the degree (with respect to some polarization) of the subvariety $X$. To achieve this, we prove a general gap principle on algebraic points that extends the new gap principle for curves embedded into their Jacobians, previously obtained by Dimitrov–Gao–Habegger and Kühne. Our new gap principle also implies the full uniform Bogomolov conjecture in abelian varieties.
@@ 1. Introduction
Throughout this article, $F$ is an algebraically closed field of characteristic 0; particularly relevant cases are $F=\overline{\mathbb{Q}}$ and $F=\mathbb{C}$. Furthermore, we let $A$ be an abelian variety defined over $F$ and consider an ample line bundle $L$ on $A$. The translates of abelian subvarieties by closed points in $A$ are called cosets.
@@
The novelty here is the complete removal of this very dependence on the ambient abelian variety $A$, confirming a folklore expectation that can be found for example in [DP07, Conj. 1.8]. It is known under the name of uniform Mordell–Lang conjecture because the number $N$ in (1.1) must depend on $g$, $d$ and $\rho$. For curves embedded in their Jacobian, it dates back to a question of Mazur [Maz86, top of p. 234], which has been answered affirmatively by work of Dimitrov, the first-named author, Habegger and the third-named author [DGH21, Küh21]. Readers may profit from the survey of the first-named author [Gao21] for an overview of these previous works and their implications for rational points on algebraic curves. Let us remark that before these works, the only uniform results of Mordell-Lang type were obtained by David and Philippon [DP07, Thm. 1.13] for subvarieties of self-products of an elliptic curve. It should be also noted that they give a completely explicit constant in this special case. In this regard, it is interesting to ask whether the present arguments can yield explicit upper bounds on the number of cosets similar to those of David and Philippon [DP07] or if substantial new ideas are necessary.
@@ 1.1. The Ueno locus and first reductions
(1.2) $X^{\circ}(F)\cap\Gamma\leq c(\dim A,\deg_{L}X)^{1+\mathrm{rk}\Gamma}.$
Besides reducing to Theorem 1.1, we also use a specialization argument of Masser [Mas89] to assume $F=\overline{\mathbb{Q}}$ in our main arguments. This is an essential reduction, since we will mostly work with arithmetical arguments around the theory of heights. Although Theorem 1.1 could itself be understood as a purely geometric assertion (e.g. for $F=\mathbb{C}$), we heavily rely on arithmetic tools throughout our proof.
@@ 1.2. A generalized gap principle
Our main contribution is a uniform count of small points, and we achieve this by establishing another kind of gap principle for algebraic points. In the case of curves in their Jacobians, preceding work [DGH21, Küh21] has been subsumed under such a New Gap Principle [Gao21, Thm. 4.1]. In this article, we take the same perspective and generalize it as follows. We say that a subvariety $X\subseteq A$ generates $A$ if the smallest abelian subvariety containing XXX-X is $A$. Furthermore, we let $\hat{h}_{L\otimes[-1]^{*}L}$ denote the Néron–Tate height on $A(\overline{\mathbb{Q}})$ associated with the symmetric ample line bundle $L\otimes[-1]^{*}L$.
@@
Given two positive integers $g$ and $d$, there exist positive constants $c_{1}=c_{1}(g,d)$ and $c_{2}=c_{2}(g,d)$ with the following property: For any polarized abelian variety $(A,L)$ of dimension $g$ and any irreducible closed subvariety $X\subseteq A$ that generates $A$ of degree $\deg_{L}X\leq d$, the set
@@
is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}(X^{\prime})<c_{2}$.
@@
We remark that Theorem 1.2 implies a uniform version of the Bogomolov conjecture, generalizing [Küh21, Thm. 3] in the case of curves embedded in their Jacobians. Since this result is of independent interest, we state it here explicitly.
@@
Given two positive integers $g$ and $d$, there exist positive constants $c_{3}=c_{3}(g,d)$ and $c_{3}=c_{3}(g,d)$ with the following property: For any abelian variety $(A,L)$ of dimension $g$ and any irreducible closed subvariety $X\subseteq A$ that generates $A$ of degree $\deg_{L}X\leq d$, we have
(1.4) $\left\{P\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes[-1]^{*}L}(P)\leq c_{3}\right\}<c_{4}.$
@@
An analogue of Theorem 1.3 for curves in algebraic tori was proven by Bombieri–Zannier [BZ95] (compare also [DP99, AD06]). For the case of abelian varieties considered here, Theorem 1.3 was known only in some cases [DP07, DKY20, Küh21]. Let us note that all these results except for [Küh21] have rather explicit constants, in contrast to our general Theorem 1.3 here.
After the first version of the current article appeared as a preprint, Yuan [Yua26] gave another proof of Theorem 1.2 in the case of curves embedded into their Jacobians previously considered in [DGH21, Küh21]. His approach has the advantage to work in the function field case (in positive characteristic) as well. It relies on previous work of Zhang, Cinkir, and de Jong [Zha93, Zha10, Cin11, dJ18] for lower bounds on the self-intersection numbers of the admissible canonical bundles of curves over global fields. Little is known in this direction for subvarieties of higher dimension and it seems very hard to generalize Yuan’s proof to the setting considered here. He also uses the new theory of Yuan–Zhang [YZ26] on adelic line bundles on quasi-projective varieties.
@@ 1.3. Ideas of the proof
    the height inequality of [DGH21, Thm. 1.6 and B.1], which allows a comparison of the Néron–Tate height and the height on the base variety, and
@@
More recently, Yuan and Zhang have reproven both these results using their general theory of adelic line bundles over quasi-projective varieties [YZ26].
@@
To prove uniform bounds as in Theorem 1.1, it is important to work with only finitely many families at all times. While this is automatic in the setting for curves embedded into their Jacobians (by an induction on the dimension of the subvariety), we have to carefully handle the invariants involved, in particular the polarization type. Even if one is only interested in the case of principal polarization, our inductive proof generally invokes abelian varieties of any polarization degree. We use various techniques to overcome these problems, including a classical result of Mumford and the Poincaré biextension on the universal abelian variety.
@@ Acknowledgements
The authors would like to thank Marc Hindry for the argument for Lemma 2.5. The authors would like to thank Gabriel Dill for his valuable comments on a previous version of the manuscript, especially for a correction of the argument for Lemma 3.1, the reference [Dil22, Lemma 2.4], discussions on the specialization argument and on the optimality of our formulation of Theorem 1.1. We thank Dan Abramovich and Philipp Habegger for comments on a draft of this paper and Abbey Bourdon for a relevant discussion on the consequence of our main result on rational points. We would like to thank the referee for their careful reading and helpful comments.
ZG received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n 945714). TG received funding from NSF grant DMS-1759514 and DMS-2100548. LK received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101027237. A revision of this article was supported by the National Science Foundation under Grant No. DMS-1928930, while the three authors were in residence at the SL-Math in Berkeley, California, during its semester on “Diophantine Geometry” in 2023.
@@ 2. Preliminaries on abelian varieties
In this section, we include some basic results on abelian varieties. Throughout the section, let $A$ be an abelian variety defined over $\overline{\mathbb{Q}}$ of dimension $g$, where $\overline{\mathbb{Q}}$ is an algebraic closure of $\mathbb{Q}$ in $\mathbb{C}$.
@@ 2.1. Polarizations
Let $\mathrm{Pic}(A)$ be the Picard group of $A$ and $\mathrm{Pic}^{0}(A)$ be the connected component of the identity. For $L\in\mathrm{Pic}(A)$, we denote its Chern class by $c_{1}(L)\in H^{2}(A(\mathbb{C}),\mathbb{Z})$. Let $A^{\vee}$ be the dual abelian variety of $A$. Then $\mathrm{Pic}^{0}(A)=A^{\vee}(\overline{\mathbb{Q}})$.
@@
The polarization is uniquely determined by the Chern class of the line bundle by the following lemma [Deb05, Theorem 6.10]:
@@
A polarized abelian variety $(A,L)$ is defined as an abelian variety $A$ equipped with an ample line bundle $L$ serving as a representative of the Chern class $c_{1}(L)$.
@@
    if $f\colon A^{\prime}\rightarrow A$ be an isogeny, then $\dim H^{0}(A^{\prime},f^{*}L)=\deg(f)\dim H^{0}(A,L)$;
@@
    there exists an abelian variety $A_{0}$, an ample line bundle $L_{0}$ on $A_{0}$ defining a principal polarization, and an isogeny $u_{0}\colon A\rightarrow A_{0}$ such that $L\cong u_{0}^{*}L_{0}$; moreover, $\deg(u_{0})=\deg_{L}(A)/g!$.
@@
For (i), (ii) and (iii), it suffices to prove the assertions over $\mathbb{C}$. Then (i) is [BL04, Corollary 3.2.8], (ii) is just the Riemann–Roch theorem [BL04, Section 3.6] and (iii) is [Deb05, Cororollary 6 to Proposition 6.12]. (iv) is [Mum74, pp. 234, Corollary 1 and its proof]. ∎
@@
We have $c_{1}(L^{\prime}\otimes L^{\otimes-1})=c_{1}(L^{\prime})-c_{1}(L)=0$. So $L^{\prime}\otimes L^{\otimes-1}\in\mathrm{Pic}^{0}(A)=A^{\vee}(\overline{\mathbb{Q}})$. The homomorphism $\phi_{L}\colon A\rightarrow A^{\vee}$, $a\mapsto t_{a}^{*}L\otimes L^{\otimes-1}$, is an isogeny because $L$ is ample. So there exists $a\in A(\overline{\mathbb{Q}})$ such that $\phi_{L}(a)=L^{\prime}\otimes L^{\otimes-1}$. Thus $L^{\prime}\cong t_{a}^{*}L$. Therefore for each $x\in A(\overline{\mathbb{Q}})$, we have
@@
Taking $x=0$, $x\in A(\overline{\mathbb{Q}})_{\mathrm{tor}}$, we get $\hat{h}_{L}(a)=0$ for the symmetric ample line bundles $L$ and $L^{\prime}$. But then $a\in A(\overline{\mathbb{Q}})_{\mathrm{tor}}$ since $L$ is symmetric ample, and hence (2.3) yields $\hat{h}_{L^{\prime}}(x)=\hat{h}_{L}(a+x)=\hat{h}_{L}(x)$ for all $x\in A(\overline{\mathbb{Q}})$. ∎
@@ 2.2. Degree estimates
$=\sum_{i=0}^{d+d^{\prime}}\binom{d+d^{\prime}}{i}(c_{1}(p_{1}^{*}L)^{i}\cap[Y])\times(c_{1}(p_{2}^{*}L)^{d+d^{\prime}-i}\cap[Y^{\prime}]).$
For reasons of dimension, the only non-vanishing term in this sum is $i=d$. So
@@
$\deg_{\alpha^{*}L^{\boxtimes 2}}(Y\times Y^{\prime})=2^{d+d^{\prime}}\binom{d+d^{\prime}}{d}\deg_{L}Y\cdot\deg_{L}Y^{\prime}\leq 4^{d+d^{\prime}}\deg_{L}Y\cdot\deg_{L}Y^{\prime}.$
@@
Let $X$ be an irreducible subvariety of $A$ and let $A^{\prime}$ denote the abelian subvariety generated by XXX-X. Then $\deg_{L}A^{\prime}\ll_{g}\deg_{L}(X)^{2g}$.
@@
$A^{\prime}=\underbrace{(X-X)+(X-X)+\cdots+(X-X)}_{\text{$g$ copies}}$
so the chain certainly becomes stationary after at most $g$ steps.
@@
$\deg_{L}(\underbrace{(X-X)+\cdots+(X-X)}_{\text{$k$ copies}})\leq 8^{2^{k}r}\deg_{L}(X)^{2g}$
@@ 3. Hilbert schemes and non-degeneracy
Let $r\geq 1$ and $d\geq 1$ be two integers. In this section, we work over $\overline{\mathbb{Q}}$, i.e. all objects and morphisms are defined over $\overline{\mathbb{Q}}$ unless stated otherwise.
Our goal is to introduce the restricted Hilbert schemes (3.6) and prove a non-degeneracy result, Proposition 3.4. This is one of the main new ingredients to prove the uniform Mordell–Lang Conjecture for higher dimensional subvarieties of abelian varieties compared to the case of curves.
@@
3.1. Parameterizing subvarieties of an abelian variety